User name peter
Log entry time 01:01:30 on June12,2002
Entry number 85903
keyword=update on right dipole
We slowly pushed up the momentum setting for the right arm, until 2.9 GeV/c at which point the current in the right dipole rose to 1070 A and eventually settled at 1036 A. We decided not to risk tripping the right dipole again by trying to push the momentum setting any higher than this.
Eric Voutier, Paul Ulmer and Werner Boeglin discussed the problem and decided to skip the Q2_a35 kinematic setting and move on the the Q2_b30 kinematic setting to take data. At this new kinematic setting, the momentum in the right arm is 2.8974 GeV/c and the current in the right dipole has settled at 1035.9 A. We have also power cycled the Q2 and Q3 in both arms, since this was not done previously when the momentum was increased just before the right dipole tripped.